bellarina:

At a large music festival in midsummer, the organizer would like to offer visitors a special bungee jumping event: when jumping on the bungee rope from a height of 48 meters, the jumper should immerse his whole body in a lake and then back off being thrown up by the rope. So that this jump does not lead to injuries, the organizer has legally secured himself and found out that the jumper must never hit the water with a speed higher than 72 over 11 m/s The function f(x) = 12 over 11 x^3 minus 81 over 11 x^2 plus 48 describes this bungee jump

itsme:

f(x)=12/11x^3-81/11x^2+48

itsme:

@supie

itsme:

I'm re-writing "f(x) = 12 over 11 x^3 minus 81 over 11 x^2 plus 48"

supie:

So I believe it is written like this; \(f(x)=12/11x^3-81/11x^2+48\) Or like this \(f(x)=\frac{12}{11}x^3-\frac{81}{11}x^2+48\)

itsme:

Yes ty ty

itsme:

does it look like this ?

itsme:

@bellarina ^